Circular No. 8881

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 T6 (CATALINA) An apparently asteroidal object discovered by the Catalina Sky Survey

(discovery observation tabulated below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been found to show cometary appearance by J. W. Young on his CCD frames taken on Oct. 14.5 UT with the Table Mountain 0.61-m f/16 Cassegrain reflector, which show a slightly elongated coma, 8′′ × 12′′ in diameter, with a short, faint 20′′ tail in p.a. 310^{o} (the coma being elongated perpendicular to the p.a. of the tail, more visible in short exposures). Young’s exposures from Oct. 15.5 show the elongated, bright coma (mag 17.2), with size and orientation of elongation the same as seen on the previous night, but now with no hint of tail.

2007 UT  α_{2000}  δ_{2000}  Mag.
Oct. 13.47047  8^{h}22^{m}02.^{s}66  +17^{o}42′00′′.0  17.8

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-T115.

T = 2007 Aug. 24.944 TT  ω = 336.^{o}865
Ω = 104.663 }2000.0
q = 2.38346 AU  i = 21.472

COMET C/2007 F1 (LONEOS) Visual total-magnitude estimates made with the comet generally < 10^{o}

above the horizon: Sept. 20.20 UT, 9.3 (J. J. Gonzalez, Leon, Spain, 0.20-m reflector); 23.20, 8.7 (Gonzalez, 25×100 binoculars); Oct. 5.20, 6.9 (Gonzalez, 10×50 binoculars); 6.16, 7.2 (B. H. Granslo, Fjellhamar, Norway, 0.10-m refractor); 7.16, 6.7 (M. Meyer, Ahlbach, Germany, 20×100 binoculars); 9.83, 7.0 (Y. Nagai, Gunma, Japan, 20×100 binoculars); 13.75, 6.3 (Meyer, Eppenrod, Germany); 14.80, 6.5 (Gonzalez).

COMET C/2007 T1 (McNAUGHT) Visual total-magnitude and coma-diameter estimates: Oct. 10.09 UT,

11.7, 2′.5 (A. Hale, Cloudcroft, NM, U.S.A. 0.41-m reflector); 12.43, 11.3, 3′.0 (M. Mattiazzo, Castlemaine, Victoria, Australia, 0.28-m reflector); 13.81, 10.4, 3′ (J. Carvajal, Avila, Spain, 0.45-m reflector); 14.82, 10.3, 4′ (J. J. Gonzalez, Leon, Spain, 25×100 binoculars).

2007 October 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green